# Thomason condition for regular algebraic stacks

### v2

@@ Abstract.
We show that the Thomason condition holds for any concentrated regular Noetherian algebraic stack with quasi-finite and locally separated diagonal. As a consequence, the derived category of quasi-coherent sheaves on such an algebraic stack is singly compactly generated. This extends results of Hall–Rydh from separated to locally separated diagonal, under additional regularity hypotheses.
@@ 1.1. What is known
Compact generation of derived categories of quasi-coherent sheaves on algebraic stacks plays a central role in modern algebraic geometry. In many cases of interest, the triangulated category $D_{\operatorname{qc}}$ is singly compactly generated. That is, there exists a single compact object whose shifts, cones, and small coproducts generate $D_{\operatorname{qc}}$.
For quasi-compact quasi-separated schemes, $D_{\operatorname{qc}}$ being singly compactly generated was established by Bondal and Van den Bergh [BV03, Theorem 3.1.1]. Extensions to various classes of Deligne–Mumford and algebraic stacks—typically in characteristic zero or under resolution property assumptions—were obtained in [Toë12, Corollary 5.2], [Kri09, proof of Proposition 5.5], and [BZFN10, §3.3].
A recent advance was made by Hall and Rydh, who proved that $D_{\operatorname{qc}}$ is singly compactly generated for any quasi-compact quasi-separated algebraic stack with quasi-finite and separated diagonal [HR17a, Theorem A]. In that work, Hall and Rydh introduced the $\beta$-Thomason condition, where $\beta$ is a cardinal. This requires not only compact generation of $D_{\operatorname{qc}}$ by a collection of cardinality $\leq\beta$, but also the existence of perfect complexes with prescribed support. Note that the $1$-Thomason condition is stronger than $D_{\operatorname{qc}}$ being singly compactly generated.
1.2. Why it matters
Before stating our results, we briefly explain the importance of the existence of compact generators for $D_{\operatorname{qc}}$. Philosophically, $D_{\operatorname{qc}}$ and its subcategories encode homological information that is expected to reflect the geometry of the underlying algebraic stack. In particular, a substantial body of work has been developed around compact objects (see e.g. [Nee92, Nee96]). This provides a manageable framework for extracting and organizing such information. These ideas have played a central role in resolving conjectures [HR17b, Nee21, Nee24] and in the study of singularity theory for schemes (see e.g. [LV26]). They have also been essential in classification results for triangulated and derived categories (see e.g. [Hal16, Lan25]) and in $K$-theoretic applications (see e.g. [Jos02a, Jos02b, Toe99, Kri09]).
1.3. What is done
The main result of this paper is the following.
Theorem 1.1 (see Corollary˜4.7).
@@ 1.3. What is done
Let $\mathcal{X}$ be a concentrated regular algebraic stack with quasi-finite and locally separated diagonal. Then $\mathcal{X}$ satisfies the $1$-Thomason condition.
Particularly, Theorem˜1.1 asserts that $D_{\operatorname{qc}}(\mathcal{X})$ is singly compactly generated, and every closed subset of $|\mathcal{X}|$ arises as the support of a perfect complex. Moreover, Theorem˜1.1 improves [Hal22, Theorem 2.1] to single compact generation in the special case of quasi-finite and locally separated diagonal. Furthermore, we prove a more general statement where $\mathcal{X}$ does not need to be concentrated; see Theorem˜4.6. Here, ‘concentrated’ is a mild condition ensuring that compact objects coincide with perfect complexes. In our context, concentratedness coincides with being ‘tame’ in the sense of [Hal16]; see e.g. [DLM25, Proposition A.1].
In contrast to [HR17a], our proof avoids étale dévissage. Instead, the proof of Theorem˜1.1 proceeds in two steps. First, using finite flat presentations of Rydh [Ryd11], we reduce to descent along an étale morphism that is finite over a closed substack. Second, we apply recollement arguments for triangulated categories to glue compact generators; see Proposition˜3.1 for a construction.
Our result extends [HR17a, Theorem A] from quasi-finite separated diagonal to quasi-finite locally separated diagonal under regularity and concentratedness hypotheses. This yields genuinely new cases even among tame Deligne–Mumford stacks, notably allowing non-separated diagonal and arbitrary characteristic.
As a consequence, for such algebraic stacks $\mathcal{X}$ in Theorem˜1.1, the natural functor $D(\operatorname{Qcoh}(\mathcal{X}))\to D_{\operatorname{qc}}(\mathcal{X})$ is an equivalence (see [HNR19, Theorem 1.2]). Furthermore, for such algebraic stacks $\mathcal{X}$ in Theorem˜1.1, the underlying topological space $|\mathcal{X}|$ determines that of the associated Balmer spectra of $D_{\operatorname{qc}}(\mathcal{X})$ (in the sense of [Bal05]); see [HR17b, Theorem C].
On a separate note, we show that the $\beta$-Thomason condition descends along a finite, flat, surjective, morphism between Noetherian algebraic stacks. See Proposition˜A.1. We record the statement mainly for its independent interest and the simplicity of the argument. In fact, [HR17a, Theorem 6.6] proves a similar statement for the case of $\beta$-crispness.
1.4. What is next
Hall and Rydh introduced the notion of $\beta$-crispness in [HR17a]. This notion requires that the Thomason $\beta$-Thomason condition be stable ascending under étale, separated, quasi-compact morphisms representable by algebraic spaces. It is not clear whether $\mathcal{S}$ in Theorem˜1.1 is $1$-crisp. In particular, the inductive argument used in the proof of Theorem˜1.1 does not appear to interact well with immersions in the sense required for $\beta$-crispness. To be precise, any étale cover of $\mathcal{S}$ need not arise as an étale cover of $\mathcal{X}$.
@@ Acknowledgments.
Lank was supported by ERC Advanced Grant 101095900-TriCatApp. The author greatly appreciates discussions with Timothy De Deyn, Jack Hall, Michal Hrbek, Kabeer Manali-Rahul, Fei Peng, and David Rydh.
@@ 2. Algebraic stacks
Our conventions for algebraic stacks are those of [Sta25]. For the derived pullback and pushforward adjunction, we adopt the conventions of [HR17a, §1] and [Ols07, LO08a, LO08b]. Unless otherwise specified, symbols such as $X$, $Y$, etc. denote schemes or algebraic spaces, while $\mathcal{X}$, $\mathcal{Y}$, etc. denote algebraic stacks. In this section, let $\mathcal{X}$ be a quasi-compact and quasi-separated algebraic stack.
@@
We specify the triangulated categories that appear in our work. Let $\operatorname{Mod}(\mathcal{X})$ denote the Grothendieck abelian category of sheaves of $\mathcal{O}_{\mathcal{X}}$-modules on the lisse-étale site of $\mathcal{X}$. Define $\operatorname{Qcoh}(\mathcal{X})$ to be the strictly full subcategory (i.e. full and closed under isomorphisms) of $\operatorname{Mod}(\mathcal{X})$ consisting of quasi-coherent sheaves. Set $D(\mathcal{X}):=D(\operatorname{Mod}(\mathcal{X}))$ for the derived category of $\operatorname{Mod}(\mathcal{X})$. Denote by $D_{\operatorname{qc}}(\mathcal{X})$ the full subcategory of $D(\mathcal{X})$ consisting of complexes with quasi-coherent cohomology sheaves. Finally, let $\operatorname{Perf}(\mathcal{X})$ denote the full subcategory of perfect complexes in $D_{\operatorname{qc}}(\mathcal{X})$.
@@
Let $M\in\operatorname{Qcoh}(\mathcal{X})$. Set $\operatorname{supp}(M):=p\big(\operatorname{supp}(p^{\ast}M)\big)\subseteq|\mathcal{X}|$ where $p\colon U\to\mathcal{X}$ is any smooth surjective morphism from a scheme. One checks that this definition is independent of the choice of $p$. Now, given any $E\in D_{\operatorname{qc}}(\mathcal{X})$, let
@@
A quasi-compact quasi-separated morphism of algebraic stacks is called concentrated if for every base change along a quasi-compact quasi-separated morphism, the derived pushforward has finite cohomological dimension. For instance, by [HR17a, Lemma 2.5(3)], morphisms which are representable by algebraic spaces are concentrated. An algebraic stack is concentrated if it is quasi-compact quasi-separated, and its structure morphism to $\operatorname{Spec}(\mathbb{Z})$ is concentrated. In fact, a quasi-compact quasi-separated algebraic stack $\mathcal{X}$ is concentrated if, and only if, any (hence all) of the following equivalent conditions hold: $\mathbf{R}\Gamma\colon D_{\operatorname{qc}}(\mathcal{X})\to D_{\operatorname{qc}}(\operatorname{Spec}(\mathbb{Z}))$ commutes with small coproducts; $\operatorname{Perf}(\mathcal{X})=D_{\operatorname{qc}}(\mathcal{X})^{c}$; or $\mathcal{O}_{\mathcal{X}}\in D_{\operatorname{qc}}(\mathcal{X})^{c}$. See [HR17a, §2, Lemma 2.5(5), & Remark 4.6] for details.
@@
Perfect complexes may be defined on any ringed site [Sta25, Tag 08G4], in particular on the lisse-étale site of $\mathcal{X}$. A complex is strictly perfect if it is a bounded complex whose terms are direct summands of finite free modules; it is perfect if it is locally strictly perfect. Let $\operatorname{Perf}(\mathcal{X})$ denote the triangulated subcategory of $D_{\operatorname{qc}}(\mathcal{X})$ consisting of perfect complexes. In general, the compact objects of $D_{\operatorname{qc}}(\mathcal{X})$ are perfect complexes [HR17a, Lemma 4.4], although the converse need not hold. The two notions coincide precisely when the algebraic stack $\mathcal{X}$ is concentrated. Any compact object of $D_{\operatorname{qc}}(\mathcal{X})$ is a perfect complex and the support of a perfect complex has quasi-compact complement (see [HR17a, Lemmas 4.4 & 4.8]).
@@
In general, $D_{\operatorname{qc}}(\mathcal{X})$ need not be compactly generated (for instance, this fails for $D_{\operatorname{qc}}(B_{k}\mathbb{G}_{a})$ when $k$ is a field of positive characteristic; see [HNR19, Proposition 3.1]). A related notion is the ‘Thomason condition’ which was introduced in [HR17a]. We say that $\mathcal{X}$ satisfies the $\beta$-Thomason condition, for some cardinal $\beta$, if $D_{\operatorname{qc}}(\mathcal{X})$ is compactly generated by a set of cardinality at most $\beta$, and if for every closed subset $Z\subseteq|\mathcal{X}|$ with quasi-compact complement there exists a perfect complex $P\in\operatorname{Perf}(\mathcal{X})$ such that $\operatorname{supp}(P)=Z$. For example, any quasi-compact quasi-separated algebraic spaces satisfies the $1$-Thomason condition (see e.g. [Sta25, Tag 08HP]).
Approximation by compacts
Let $(T,E,m)$ be a triple consisting of a closed subset $T\subseteq|\mathcal{X}|$, $E\in D_{\operatorname{qc}}(\mathcal{X})$, and $m\in\mathbb{Z}$. We say that approximation by compacts holds for $(T,E,m)$ if there exists a $P\in D_{\operatorname{qc}}(\mathcal{X})^{c}$ with support contained in $T$ and a morphism $P\to E$ such $\mathcal{H}^{i}(P)\to\mathcal{H}^{i}(E)$ is an isomorphism for $i>m$ and surjective when $i=m$. More generally, we say $\mathcal{X}$ satisfies approximation by compacts if for every closed subset $T\subseteq|\mathcal{X}|$ with $\mathcal{X}\setminus T\to\mathcal{X}$ quasi-compact there exists an integer $r$ such that for any triple $(T,E,m)$ with $E$ is (mr)(m-r)-pseudocoherent (see e.g. [Sta25, Tag 08FT]) and $\mathcal{H}^{i}(E)$ is supported on $T$ for $i\geq m-r$, approximation by compacts holds. See [HLLP25, $\S 3$] for details. For example, any quasi-compact quasi-separated algebraic stack with quasi-finite and separated diagonal satisfies approximation by compacts [HLLP25, Corollary 5.4].
@@ 3. Recollements
We briefly recall the notion of a recollement. See [BBDG18, §1.4] for details. A recollement is a commutative diagram of triangulated categories and exact functors of the form
@@
$D_{\operatorname{qc}}(\mathcal{U})$$D_{\operatorname{qc}}(\mathcal{X})$$D_{\operatorname{qc},Z}(\mathcal{X})$$\scriptstyle{\mathbf{R}j_{\ast}}$$\scriptstyle{\mathbf{L}j^{\ast}}$$\scriptstyle{j^{!}}$$\scriptstyle{i^{!}}$$\scriptstyle{i_{\ast}}$$\scriptstyle{i^{\ast}}$
where in particular $i_{\ast}$ is the natural inclusion and $j^{!}$ the right adjoint of $\mathbf{R}j_{\ast}$.
@@
First, we spell out the existence of the needed functors. By [HR17a, Theorem 4.14(1)], the right adjoint of $\mathbf{R}j_{\ast}$ exists, which we denote by $j^{!}$. Now, there is a Verdier localization
@@
This can be seen using e.g. [HR17b, Example 1.2]. Note that $\mathbf{R}j_{\ast}$ is right adjoint to $\mathbf{L}j^{\ast}$ on $D_{\operatorname{qc}}$ (see e.g. [HR17a, §1.3]). So, by [GP18, Lemma 2.2(ii)], $i_{\ast}$ must admit a right adjoint as well, which we denote by $i^{!}$ (loc. cit. pulls from [CPS88b, Theorem 1.1] and [CPS88a, Theorem 2.1]). It follows that
@@
is a Verdier localization sequence. However, being that $j^{!}$ is the right adjoint of $\mathbf{R}j_{\ast}$ on $D_{\operatorname{qc}}$, we can apply [GP18, Lemma 2.2(ii)] once more to see that $i^{!}$ admits a right adjoint as well, which we denote by $i^{\ast}$. Tying things together, we have the required data for a recollement; i.e. a Verdier localization sequence which is a localization and colocalization sequence in the sense of [Kra10, §4]. ∎

### v3

# Perfect generation for regular algebraic stacks

@@ Abstract.
We show that the derived category of complexes with quasi-coherent cohomology on a regular Noetherian algebraic stack with quasi-finite diagonal is generated by a single perfect complex. In the concentrated case, the category is singly compactly generated. Key ingredients in the proofs include gluing generators along recollement and the use of suitable filtrations and presentations of the algebraic stack.
@@ 1.1. What is known
Consider the derived category $D_{\operatorname{qc}}$ of complexes with quasi-coherent cohomology on an algebraic stack. It is classical that $D_{\operatorname{qc}}$ is singly compactly generated for quasi-compact quasi-separated schemes [BV03, Theorem 3.1.1]. However, the situation is different for algebraic stacks, e.g. perfect complexes $P\in D_{\operatorname{qc}}$ need not be compact.
Early extensions for single object compact generation to various classes of algebraic stacks were obtained in [TOË12, Corollary 5.2], [KRI09, proof of Proposition 5.5], and [BFN10, §3.3]. More recently, the work of Hall–Rydh has pushed these results further. These include quasi-compact quasi-separated algebraic stacks with quasi-finite separated diagonal [HR17a, Theorem A] and quasi-compact quasi-separated Deligne–Mumford $\mathbb{Q}$-stacks [HR18, Theorem 7.4].
It was shown in [HAL22] that $D_{\operatorname{qc}}$ is compactly generated for any affine-pointed concentrated regular algebraic stack of finite Krull dimension. However, it is not clear whether single object compact generation occurs. Here, ‘concentrated’ is a mild condition ensuring that compact objects coincide with perfect complexes. In the case of finitely presented inertia, being concentrated wth quasi-finite diagonal coincides with being ‘tame’ in the sense of [HAL16]; see e.g. [DLM25b, Proposition A.1]. Also, ‘affine-pointed’ means that every morphism from the spectrum of a field is affine.
1.2. What is new
Our focus is on single object compact generation for regular algebraic stacks. As a first step, we consider a weaker notion of generation in triangulated categories. Recall that a collection $\mathcal{B}$ in a category $\mathcal{C}$ is said to generate $\mathcal{C}$ if for any nonzero $E\in\mathcal{C}$ there is an $n\in\mathbb{Z}$ and $B\in\mathcal{B}$ such that $\operatorname{Hom}(B[n],E)\not\cong 0$.
We establish generation by a single perfect complex.
Theorem 1.1.
@@ 1.2. What is new
Let $\mathcal{X}$ be a quasi-compact quasi-separated regular algebraic stack with quasi-finite diagonal. Then there is a $P\in\operatorname{Perf}(\mathcal{X})$ which generates $D_{\operatorname{qc}}(\mathcal{X})$.
Note that we do not assume separated diagonal or that $\mathcal{X}$ is concentrated. Moreover, Theorem˜1.1 yields genuinely new cases, even among Deligne–Mumford stacks in arbitrary characteristic under the regularity hypothesis (e.g. smooth over a field or DVR). It appears that generation by a single perfect complex which need not be compact seems quite new.
The proof proceeds in two steps. First, we use the presentations of [HR18] for algebraic stacks with finite stabilizers. This allows us to argue inductively on the length of a monomorphic splitting sequence of a Nisnevich covering. Second, we use recollement techniques for triangulated categories to glue generators (see Proposition˜3.1). We find the use of recollements here to be of independent interest.
This reduces the problem to detecting single object generation on suitable locally closed substacks. We then apply finite duality to show that subcategories of the form $D_{\operatorname{qc},Z}$ are generated by a single object, where the regularity hypothesis is used. In fact, the argument suggests single object compact generation, but we are not able to prove this (see Remark˜4.5).
A straightforward application is the following.
Corollary 1.2.
$D_{\operatorname{qc}}(\mathcal{X})$ is singly compactly generated for any concentrated regular algebraic stack $\mathcal{X}$ with quasi-finite diagonal.
Under the quasi-finite diagonal condition, this result improves [HAL22, Theorem 2.1] to single object compact generation and removes finiteness of Krull dimension. It applies to the case $\mathcal{X}$ is a regular Noetherian quasi-DM $\mathbb{Q}$-stack with affine stabilizers because concentratedness follows from [HR15, Theorem 2.1(1)]. Furthermore, it applies when $\mathcal{X}$ is smooth, finitely presented, and quasi-DM over a DVR (possibly of mixed characteristic) where the stabilizers of $\mathcal{X}$ are affine and nice (in the sense of [HR15]).
@@ Acknowledgments.
Lank was supported by ERC Advanced Grant 101095900-TriCatApp. The author greatly appreciates comments and/or discussions with Jarod Alper, Timothy De Deyn, Jack Hall, Michal Hrbek, Kabeer Manali-Rahul, Fei Peng, and David Rydh.
@@ 2. Algebraic stacks
Our conventions for algebraic stacks are those of [STA26]. For the derived pullback and pushforward adjunction, we adopt the conventions of [HR17a, §1] and [OLS07, LO08a, LO08b]. Unless otherwise specified, symbols such as $X$, $Y$, etc. denote schemes or algebraic spaces, while $\mathcal{X}$, $\mathcal{Y}$, etc. denote algebraic stacks. In this section, let $\mathcal{X}$ be a quasi-compact quasi-separated algebraic stack.
@@
We specify the triangulated categories that appear in our work. Let $\operatorname{Mod}(\mathcal{X})$ denote the Grothendieck abelian category of sheaves of $\mathcal{O}_{\mathcal{X}}$-modules on the lisse-étale site of $\mathcal{X}$. Define $\operatorname{Qcoh}(\mathcal{X})$ to be the strictly full subcategory (i.e. full and closed under isomorphisms) of $\operatorname{Mod}(\mathcal{X})$ consisting of quasi-coherent sheaves. Set $D(\mathcal{X}):=D(\operatorname{Mod}(\mathcal{X}))$ for the derived category of $\operatorname{Mod}(\mathcal{X})$. Denote by $D_{\operatorname{qc}}(\mathcal{X})$ the full subcategory of $D(\mathcal{X})$ consisting of complexes with quasi-coherent cohomology. Finally, let $\operatorname{Perf}(\mathcal{X})$ denote the full subcategory of perfect complexes in $D_{\operatorname{qc}}(\mathcal{X})$.
@@
Let $M\in\operatorname{Qcoh}(\mathcal{X})$. Set $\operatorname{supp}(M):=p\big(\operatorname{supp}(p^{\ast}M)\big)\subseteq|\mathcal{X}|$ where $p\colon U\to\mathcal{X}$ is any smooth surjective morphism from a scheme. One can check that this definition is independent of the choice of $p$. Now, given any $E\in D_{\operatorname{qc}}(\mathcal{X})$, let
@@
We record two lemmas for convenience.
Lemma 2.1.
Let $f\colon\mathcal{Y}\to\mathcal{X}$ be a morphism of quasi-compact quasi-separated algebraic stacks. Consider a closed $Z\subseteq|\mathcal{X}|$. Then
Proof.
Hence, we can reduce to the case of affine schemes. In this setting, $D_{\operatorname{qc},Z}(\mathcal{X})$ is compactly generated (see [ROU08, Theorem 6.8]). From Remark˜3.2, we know that $D_{\operatorname{qc},Z}(\mathcal{X})$ is compactly generated by a subset $\mathcal{B}$ of $D_{\operatorname{qc},Z}(\mathcal{X})\cap\operatorname{Perf}(\mathcal{X})$. Furthermore, every object of $D_{\operatorname{qc},Z}(\mathcal{X})$ is a homotopy colimit of iterated extensions of small coproducts of shifts of objects in $\mathcal{B}$ (see [STA26, Tag 09SN]). Therefore, the desired claim can be checked if we can show $\mathbf{L}f^{\ast}\mathcal{B}\subseteq D_{\operatorname{qc},f^{-1}(Z)}(\mathcal{Y})$. However, [HR17a, Lemma 4.8(2)] tells us
which completes the proof. ∎
Lemma 2.2.
Let $f\colon\mathcal{Y}\to\mathcal{X}$ be a proper morphism of Noetherian algebraic stacks which is representable by scheme. For any closed subset $Z\subseteq|\mathcal{Y}|$, one has $\mathbf{R}f_{\ast}D_{\operatorname{qc},Z}(\mathcal{Y})\subseteq D_{\operatorname{qc},f(Z)}(\mathcal{X})$.
Proof.
Here, $f$ is concentrated. So, the problem is smooth local via [HR17a, Theorem 2.6(2)]. Hence, we can impose Noetherian schemes $X$ and $Y$. In such a case, $D_{\operatorname{qc},Z}(Y)$ is compactly generated by perfect complexes on $Y$ which have support contained in $Z$ (see [ROU08, Theorem 6.8]). Furthermore, every object of $D_{\operatorname{qc},Z}(Y)$ is a homotopy colimit of iterated extensions of small coproducts of shifts of objects in $\mathcal{B}$ (see [STA26, Tag 09SN]). Therefore, the desired claim can be checked if we can show $\mathbf{R}f_{\ast}\mathcal{B}\subseteq D_{\operatorname{qc},f(Z)}(X)$. However, this is known for proper morphisms of Noetherian schemes, see e.g. [GW23, Remark 23.46(2)]. ∎
@@
A quasi-compact quasi-separated morphism of algebraic stacks is called concentrated if for every base change along a quasi-compact quasi-separated morphism, the derived pushforward has finite cohomological dimension. For instance, by [HR17a, Lemma 2.5(3)], morphisms which are representable by algebraic spaces are concentrated. An algebraic stack is concentrated if it is quasi-compact quasi-separated, and its structure morphism to $\operatorname{Spec}(\mathbb{Z})$ is concentrated. In fact, a quasi-compact quasi-separated algebraic stack $\mathcal{X}$ is concentrated if, and only if, $\operatorname{Perf}(\mathcal{X})=D_{\operatorname{qc}}(\mathcal{X})^{c}$ if, and only if, $\mathcal{O}_{\mathcal{X}}\in D_{\operatorname{qc}}(\mathcal{X})^{c}$. See [HR17a, §2, Lemma 2.5(5), & Remark 4.6] for details.
@@
Perfect complexes may be defined on any ringed site [STA26, Tag 08G4], in particular on the lisse-étale site of $\mathcal{X}$. A complex is strictly perfect if it is a bounded complex whose terms are direct summands of finite free modules; it is perfect if it is locally strictly perfect. Let $\operatorname{Perf}(\mathcal{X})$ denote the triangulated subcategory of $D_{\operatorname{qc}}(\mathcal{X})$ consisting of perfect complexes. In general, the compact objects of $D_{\operatorname{qc}}(\mathcal{X})$ are perfect complexes [HR17a, Lemma 4.4], although the converse need not hold. The two notions coincide precisely when the algebraic stack $\mathcal{X}$ is concentrated. Any compact object of $D_{\operatorname{qc}}(\mathcal{X})$ is a perfect complex and the support of a perfect complex has quasi-compact complement (see [HR17a, Lemmas 4.4 & 4.8]).
@@
In general, $D_{\operatorname{qc}}(\mathcal{X})$ need not be compactly generated (for instance, this fails for $D_{\operatorname{qc}}(B_{k}\mathbb{G}_{a})$ when $k$ is a field of positive characteristic; see [HNR19, Proposition 3.1]). A related notion is the ‘Thomason condition’ which was introduced in [HR17a]. We say that $\mathcal{X}$ satisfies the $\beta$-Thomason condition, for some cardinal $\beta$, if $D_{\operatorname{qc}}(\mathcal{X})$ is compactly generated by a set of cardinality at most $\beta$, and if for every closed subset $Z\subseteq|\mathcal{X}|$ with quasi-compact complement there exists a perfect complex $P\in\operatorname{Perf}(\mathcal{X})$ such that $\operatorname{supp}(P)=Z$. For example, any quasi-compact quasi-separated algebraic spaces satisfies the $1$-Thomason condition (see e.g. [STA26, Tag 08HP]).
@@ 3. Recollements
3.1. Reminder
We briefly recall the notion of a recollement. See [BBD+18, §1.4] for details. A recollement is a commutative diagram of triangulated categories and exact functors of the form
@@ 3.1. Reminder
3.2. A special case
We record some results in setting of algebraic stacks. The following is well-known for schemes (see e.g. [JØR09, Theorem 1]).
@@ 3.2. A special case
$D_{\operatorname{qc}}(\mathcal{U})$$D_{\operatorname{qc}}(\mathcal{X})$$D_{\operatorname{qc},Z}(\mathcal{X})$$\scriptstyle{\mathbf{R}j_{\ast}}$$\scriptstyle{\mathbf{L}j^{\ast}}$$\scriptstyle{j^{\times}}$$\scriptstyle{i^{!}}$$\scriptstyle{i_{\ast}}$$\scriptstyle{i^{\ast}}$
where $i_{\ast}$ is the natural inclusion and $j^{\times}$ the right adjoint of $\mathbf{R}j_{\ast}$.
@@
First, we spell out the existence of the needed functors. Recall that the right adjoint $j^{\times}$ of $\mathbf{R}j_{\ast}$ exists because $j$ is concentrated (see [HR17a, Theorem 4.14(1)]). By [HR17b, Example 1.2], there is a Verdier localization
@@
Note that $\mathbf{R}j_{\ast}$ is right adjoint to $\mathbf{L}j^{\ast}$ on $D_{\operatorname{qc}}$ (see [HR17a, §1.3]). So, by [GP18, Lemma 2.2(ii)], $i_{\ast}$ must admit a right adjoint, which we denote by $i^{!}$ (loc. cit. pulls from [CPS88b, Theorem 1.1] and [CPS88a, Theorem 2.1]). It follows that
@@
is a Verdier localization sequence. However, being that $j^{\times}$ is the right adjoint of $\mathbf{R}j_{\ast}$ on $D_{\operatorname{qc}}$, we can apply [GP18, Lemma 2.2(ii)] once more to see that $i^{!}$ admits a right adjoint as well, which we denote by $i^{\ast}$. Tying things together, we have the required data for a recollement; i.e. a Verdier localization sequence which is a localization and colocalization sequence in the sense of [KRA10, §4]. ∎